OpenStudy (anonymous):

Create a function, consisting of two different operations. Prove that your function is a legitimate function. Using your function, explain to how to solve for f(3). Show your work and explain each step using complete sentences. Also, explain how to find the inverse of your function

OpenStudy (zzr0ck3r):

write down any math statement with addition and multiplication

OpenStudy (zzr0ck3r):

then put f(x) = in front of it

OpenStudy (anonymous):

f(x)=6x-9=4x+7

OpenStudy (zzr0ck3r):

well only one =

OpenStudy (zzr0ck3r):

so f(x) = 6x-9

OpenStudy (zzr0ck3r):

this is a function from R to R, you can prove its a function by checking two properties 1) f(a) is defined for all a in R...this is true because you can put any real number as an input to this function 2) a=b implies f(a) = f(b) you can prove this by starting with the assumption a=b then multiply both sides by 6 6a=6b then subtract 9 from both sides 6a-9 = 6b-9 in other words f(a) = f(b) so our "thing" is indeed a function. can you find f(3)?

OpenStudy (anonymous):

Ok, but I need two different operations

OpenStudy (zzr0ck3r):

6*x+(-9)

OpenStudy (zzr0ck3r):

that is two

OpenStudy (zzr0ck3r):

multiplication and addition

OpenStudy (anonymous):

No, what I mean is on operation on one side and another operation on the other side.

OpenStudy (zzr0ck3r):

that is not what its asking

OpenStudy (zzr0ck3r):

f(x) will always be by itself on one side.....

OpenStudy (zzr0ck3r):

A math statement is something WIHTOUT a = sign A math equation is something with an = sign A function is f(x) = something

OpenStudy (anonymous):

Oh okay, I understand

OpenStudy (anonymous):

So how can I prove its f(x) = 6x-9 a legitimate function?

OpenStudy (zzr0ck3r):

I showed you \(\uparrow\)

OpenStudy (anonymous):

Oh, okay. So I just plug in three for the f(3) part, and how about inverse?

OpenStudy (zzr0ck3r):

to find the inverse, switch the x, and y then solve for y \(y=6x-9 \) becomes \(x=6y-9\) solve for y

OpenStudy (zzr0ck3r):

\(y=\frac{x+9}{6}\)

OpenStudy (zzr0ck3r):

so \(f^{-1}(x) = \frac{x+9}{6}\)

OpenStudy (anonymous):

How could I solve this?

OpenStudy (zzr0ck3r):

\(x=6x-9\iff x+9=6y\iff \frac{x+9}{6}=y=f^{-1}(x)\)

OpenStudy (anonymous):

Do I have to find for either variables?

OpenStudy (zzr0ck3r):

im confused by the question

OpenStudy (anonymous):

So my function is f(x)=6x-9 Solving for f(3) is f(3)=6(3)-9 And the inverse is what you wrote above.

OpenStudy (zzr0ck3r):

correct

OpenStudy (anonymous):

For my answer I'd write exactly what you wrote?

OpenStudy (anonymous):

For the inverse?

OpenStudy (zzr0ck3r):

yes that would work, but I fear you are not understanding it all.. please ask if you have any questions. The reason I say this, is because when someone asks " so thats what I put as the answer" usually means they are not even sure what the question is asking ( what is a inverse of a function ).

OpenStudy (anonymous):